I think this is the category this problem falls under:
(15.91)
1.96 --------
("square root" 15)
I don't know what to do, I'm pretty sure I would take the square root of 15 and times it by itself and the 15.91 and then that cancels out the bottom of the fraction. Then I have 1.96 (15.91) square root of 15. Then I would take the square root of 15, well that's where I get lost. I can't take anything out of 15 I dont think. I mean, I could take 6 out leaving 9 in? I know there's something like that, that I'm supposed to do. But I'm just not sure. Please help!

Answer by checkley71(8403)

1.96*15.91/sqrt15=31.1836/sqrt15=31.1836/3.873=8.0515362 answer.